OpenStudy (anonymous):

Given that f(x) = 6x + 2 and g(x) = the quantity of 2x plus 4 divided by 5, solve for g(f(1)). 1 3 4 8

OpenStudy (anonymous):

i need some help please

OpenStudy (anonymous):

@sidona005 can you help?

OpenStudy (anonymous):

you plug 1 in for 6x+2 and your answer would be 8 so now you would plug in 8 for all the x's in g so 2(8)+4=20 now divide 20 by the 5 and you get 4

OpenStudy (gorv):

first find f(1)

OpenStudy (gorv):

f(1)=6*1+2=8

OpenStudy (gorv):

\[g(x)=\frac{ 2x+4 }{ 5 }\]

OpenStudy (gorv):

g(f(1)0=g(6)=(2*6+4)/5